OpenStudy (anonymous):

Match each proportion with its solution. Match Term Definition x over two equals fifteen over six A) forty twelve over sixteen equals thirty over x B) 6 four over twelve equals x over eighteen C) 5 twelve over x equals eight over twelve D) eighteen

OpenStudy (mertsj):

\[\frac{x}{2}=\frac{15}{6}, 6x=30, x=5\]

OpenStudy (mertsj):

\[\frac{12}{16}=\frac{30}{x}, 12x=480, x=30\]

OpenStudy (mertsj):

\[\frac{4}{12}=\frac{x}{18}, 12x=72, x=6\]

OpenStudy (mertsj):

\[\frac{12}{x}=\frac{8}{12}, 8x=144, x=18\]